OpenStudy (anonymous):

3x + 2x - 7 = 18 x =

Nnesha (nnesha):

combine like terms 3x + 2x = ???

Nnesha (nnesha):

you can combine them bec 3 and 2 both have same exponent of same variable

Nnesha (nnesha):

what you get 3x + 2x = ???

OpenStudy (anonymous):

5x

Nnesha (nnesha):

yes right now you have  now you have to move -7 so add 7 both side

OpenStudy (anonymous):

(2x+3x)=18+7 5x=25 x=25/5 x=5

Nnesha (nnesha):

 let me know what you get :)